OpenStudy (anonymous):

What is the value of the missing exponent? (Type only the numerical value of the missing exponent. Do not skip any spaces.) (2x3)?=4x6

OpenStudy (anonymous):

\[\Large (2x^3)^?=4x^6 \] Hint: 6/3=2 4/2=2 Then whats da value of the missing exponent ?